# Self-Admitted GenAI Usage in Open-Source Software

### v2

@@ 5.1 Method
The GenAI mentions we identified as part of RQ1 allow us to approximate the point in time when the open-source projects in our sample started using the GenAI tools. We included 151 projects with true positive GenAI mentions that, according to our analysis for RQ2, did not prohibit or discourage the use of GenAI tools. Hence, we use self-admitted GenAI usage as a proxy for GenAI tool adoption.
@@
Specifically, we defined code churn as the percentage of lines that are reverted or updated within 14 days after they were initially added or modified. We added a second definition that focuses on churned files instead of lines to gain a more comprehensive understanding of the impact of GenAI adoption on the selected repositories,
@@
$ch_{L}(c)=\frac{\text{\#lines changed again within 14 days}}{\text{total \#lines changed by }c}$
@@
$ch_{F}(c)=\frac{\text{\#files changed again within 14 days}}{\text{total \#files changed by }c}$
@@
    The average churn per repository, comparing pre- and post-GenAI adoption using Wilcoxon signed-rank test [23]. We applied the Wilcoxon Z statistic $r$, to measure the paired effect and interpreted the effect size as follows [24]: $|r|<0.1$ as negligible, $0.30.1\leq|r|<$ 0.3 as small, $0.50.3\leq|r|<$ 0.5 as medium, and $0.5\leq|r|$ as large;
@@
    The average churn over all commits in all repositories, again comparing pre- and post-GenAI adoption using Mann-Whitney test [25]. We applied the Cliff $\delta$ [26], to mesaure the independent effect and interpreted the effect size as follows [27]: $|\delta|<0.147$ as negligible, $0.330.147\leq|\delta|<$ 0.33 as small, $0.4740.3\leq|\delta|<$ 0.474 as medium, and $0.474\leq|\delta|$ as large.
@@
We categorized the patterns that emerged from the RDD analysis based on two key characteristics: (1) trend and (2) slope. The trend characteristic captures whether the code churn exhibits an upward or downward trend when comparing pre- and post-GenAI adoption periods. The slope characteristic captures whether and how the slope of the trend line changes before and after GenAI adoption. The Ordinary Least Squares (OLS) model used as part of RDD requires a minimum of five weeks of data to estimate the four parameters intercept, time trend, treatment effect, and interaction while maintaining positive degrees of freedom [29] to ensure that there are enough data points to estimate the model parameters without overfitting. After applying a threshold of at least one commit per week over five weeks, we had to exclude two repositories without sufficient data in the pre-GenAI adoption period. For the 149 included repositories, we identified four patterns:
@@
These patterns provide a useful framework for analyzing how code churn metrics change after GenAI adoption in different project contexts.
@@ 5.2 Results
Table VII illustrates the variations in code churn of the studied repositories. Of the 151 repositories with self-admitted GenAI usage, 119 had a significant difference in file-based churn and 120 had a significant difference in line-based churn ($p<0.05$). Eleven repositories had an increasing file-based churn with a medium-to-large effect size, and 15 had an increasing line-based-churn with a medium-to-large effect size. A decreasing churn was more common: 49 repositories had a decreasing file-based churn with a medium-to-large effect size, and 49 repositories had a decreasing line-based churn with a medium-to-large effect size.
@@
These results are contrary to our expectations because the GitClear report was very bold in claiming that code churn increased for the projects they studied, suggesting a downward pressure on code quality [11]. While we observed that some repositories have an increasing trend in code churn, both the overall trend and the trend and many individual repositories points to a decreasing code churn over time. Therefore, with our data and methodology, we cannot confirm this claim.
Table VIII summarizes the results of our RDD analysis. We observed that only 26 (file-based) respectively 30 (line-based) repositories showed significant code churn trends ($p<0.05$). For file-based churn, an overall upward trend with a negative slope after the cutoff date was most common (12 repositories). For line-based churn, an overall downward trend with a negative slope was most common (12 repositories). However, there were almost as many repositories (10) with an overall downward trend, but a positive slope after the cutoff date. Figure 1 presents examples of all four patterns that we observed, and the complete RDD results are available as part of our replication package. In summary, while we found 12 repositories with a downward trend in line-based churn and a negative slope, overall we cannot conclude that code churn is rising.
Summary RQ3: Our results revealed that for most of the repositories analyzed, there was no significant change in code churn after GenAI adoption. However, we did find 12 repositories with an overall downward trend in line-based code churn and a negative slope after the first GenAI mention. This indicates that more research is required to understand why certain projects are affected and others not and how higher (or lower) code churn relates to the long-term maintainability of software projects.
@@ 6 Discussion
In this section, we discuss and contextualize the results across our three research questions and summarize the implications for software developers and researchers.
@@ 6.1 RQ1: Reasons for Mentioning GenAI Tools
Our analysis revealed that developers primarily use GenAI tools for code generation, natural language translation, and code refactoring. Tufano et al. [6] explored mentions of ChatGPT in commits, PRs, and issues. They identified that the three most common task categories were feature implementation and enhancement, software quality, and documentation. In our study, we present a more fine-grained and comprehensive categorization of tasks automated by both ChatGPT and GitHub Copilot. For studies targeting GitHub repositories, it is crucial to consider GitHub Copilot as well, because (1) opposed to ChatGPT, it is a tool tailored to software development, and (2) it is more deeply embedded in developers’ workflows (their local editors, but also into the GitHub platform as a whole). Moreover, we complement the task categories by specifically discussing content and usage purposes. In addition, we identified patterns of human intervention. Hou et al. [1] reviewed literature on LLMs for software engineering. They found that software engineering research has a strong focus on code generation and program repair. We complement this observation with a detailed taxonomy of how open-source developers use LLM-based tools in their projects. In addition to code generation, we found that internationalization and translation of natural language are common use cases for LLMs in open-source software projects. We further found instances of projects regulating the usage of GenAI tools, which we analyzed in more detail as part of RQ2. While our study partially confirms previous studies on software development tasks being automated using GenAI tools, we contribute three novel perspectives: (1) some developers deeply care about acknowledging GenAI usage in open-source software, (2) open-source maintainers try to actively guide and regulate GenAI usage, and (3) issues with generated code can trigger human interventions in open-source software projects.
@@
For researchers, our notion of self-admitted GenAI usage, inspired by self-admitted technical debt [10], can be a valuable lens for studying GenAI usage in practice. Of course, only a fraction of the generated software artifacts contain GenAI mentions and the artifacts that are documented might not be representative of the overall GenAI usage. Better understanding when and why developers decide to self-admit GenAI usage is one potential direction for future work. Another direction is to build a classifier that automatically labels true positive GenAI mentions according to the definition presented in Section 2.2. Our annotated dataset can serve as a starting point. An improved and scaled detection of self-admitted GenAI usage would allow researchers to build larger datasets that could then enable more comprehensive studies on code quality and maintainability of generated code.
@@
Software developers can browse our taxonomy of tasks, content types, and purposes to identify potential applications of GenAI tools in their projects. One central aspect is whether to establish guidelines clarifying in which cases project maintainers require contributors to disclose and acknowledge GenAI usage (see also Section 6.2). In our study, we found instances where such acknowledgments were combined with warnings about the potential negative implications of GenAI tool usage. Sometimes, GenAI tools were also blamed for issues. However, acknowledgment can also serve a positive purpose. We found GenAI mentions in the context of documented prompts. The question arises not only when to acknowledge GenAI usage, but also which context to document beyond the tool name (which we focused on). In which cases does it make sense to document complete prompts and where and how should one document the generation context? Such questions can be addressed both from a scientific and from a more practice-oriented perspective. Our findings suggest that a more standardized approach for documenting GenAI contributions is required, since most self-admitted GenAI usages did not document the generation context beyond brief summaries.
@@ 6.2 RQ2: Existing Guidelines for GenAI Usage
Motivated by the purpose categories Documentation and Acknowledgement and Guidance and Best Practices that we identified while answering RQ1 (see Table V), we further explored the policies and usage guidelines for GenAI tools that we found (see Table VI). Their content ranged from encouraging developers to use GenAI tools to prohibiting their usage entirely. The developer survey we conducted confirmed the broad spectrum of positions, covering ethical, legal, and practical considerations.
@@
Mentioned aspects include the unclear copyright situation of the training data, the unclear implications for generated content, data privacy risks when sharing inputs with GenAI systems, and concerns regarding code quality and maintainability. Moreover, a majority of our survey participants agreed that the regulation of GenAI usage is necessary in open-source projects. Related to that, participants argued for transparent disclosure of GenAI usage and also for documenting the generation context. It is unclear how much transparency is required and what purposes it can serve in the future: Is a binary flag sufficient? Or is it better to document the percentage of generated content, as suggested by one participant? Or the whole prompt? Do only manually written prompts need to be disclosed, or also system prompts? This aspect is aligned with the questions raised in the discussion for RQ1 regarding prompt context.
@@
Our results suggest that software developers, especially those maintaining open-source software projects, should articulate a clear position regarding GenAI usage in their projects. Many positions are possible. The spectrum ranges from a general recommendation to use GenAI tools, over recommendations for specific tools and use cases, to more restrictive policies requiring an extensive peer review of generated content, or policies prohibiting GenAI usage completely. Open-source projects should clearly communicate expectations regarding GenAI usage to their contributors. For downstream consumers of open-source dependencies, explicit GenAI policies serve as a signal of due diligence that may influence their dependency selection.
@@
Our analysis of policies, guidelines, and developers’ positions regarding GenAI regulation provides a solid foundation for researchers to design and conduct further studies on how software projects regulate GenAI usage and how such regulations impact development activity. An idea worth exploring is whether existing GenAI tools could be augmented to capture provenance information during generation that could be automatically documented in source code comments, comment messages, or artifacts such as Software Bills of Materials (SBOMs) [32] or Software Bill of Materials for AI (SBOM for AI) [33]. There are already open-source projects that extensively document prompts in commit messages.222github.com/cloudflare/workers-oauth-provider/commit/adcb… This provenance information is essential for studying the long-term impact of code generation on maintainability, but also for facilitating software supply chain transparency and effective vulnerability management. Researchers can contribute to the development of standardized metadata formats for capturing provenance and tracability information for source code, but also for other software artifacts.
@@ 6.3 RQ3: Impact of GenAI Usage on Code Churn
Our results for RQ3 challenge popular narratives about the impact of GenAI on software development. Contrary to claims in the GitClear report, which was extensively discussed in the software development community,333news.ycombinator.com/item?id=39177008444reddit.com/r/…/new_github_copilot_research_finds_downward/ we did not find an increasing code churn after GenAI adoption. The overall trend we observed pointed in the opposite direction, that is, we noticed a decreasing average code churn. This is in line with a study by Grewal et al. [34] which examined how ChatGPT-generated code is integrated into GitHub projects. They found that approximately 54% of the generated code lines were integrated and only 2.5% of them were later modified. However, we did find 12 repositories for which we could confirm that churn is increasing since GenAI adoption.
Due to the contrasting evidence, researchers need to further explore the factors that contribute to increased code churn. The patterns we identified using our RDD analysis are a valuable lens for clustering projects, to guide a detailed qualitative study of projects exhibiting certain patterns. The difference between our results and the GitClear report can be partially attributed to the methodological differences between the studies. While GitClear used a global cutoff date, we used the first GenAI mention in a repository as a proxy for GenAI adoption, thus following a more fine-grained approach. Moreover, we introduced file-level and line-level code churn and analyzed data both on the project-level and globally. Our definitions and the code we share as part of our research artifact enable other researchers to consider code churn in their own studies.
For software developers, our results suggest that the impact of GenAI adoption on the development activity in software projects might not be as clear as suggested by the GitClear report. Considering that we did notice an increasing code churn in several projects, it is nevertheless important for project maintainers to monitor their development activity and the quality of contributions. Going forward, we might extend our implementation to calculate code churn into a tool that GitHub project maintainers can easily integrate into their repositories.

### v3

@@ 5.1 Method
The GenAI mentions we identified as part of RQ1 allow us to approximate the point in time when the 156 open-source projects in our sample started using the GenAI tools. We included 151 repositories with true positive GenAI mentions that did not prohibit the use of GenAI tools. That is, we excluded five repositories prohibiting GenAI usage according to our RQ2 analysis. Hence, we use self-admitted GenAI usage as a proxy for GenAI tool adoption.
@@
Specifically, we defined code churn as the percentage of lines that are reverted or updated within 14 days after they were initially added or modified. We added a second definition that focuses on churned files instead of lines to gain a more comprehensive understanding of the impact of GenAI adoption on the selected repositories.
@@
[…]
@@
[…]
@@
    The average churn per repository, comparing pre- and post-GenAI adoption using Wilcoxon signed-rank test [49]. We applied the Wilcoxon Z statistic $r$, to measure the paired effect and interpreted the effect size as follows [5]: $|r|<0.1$ as negligible, $0.1\leq|r|<0.3$ as small, $0.3\leq|r|<0.5$ as medium, and $0.5\leq|r|$ as large;
@@
    The average churn over all commits in all repositories, comparing pre- and post-GenAI adoption using Mann-Whitney test [21]. We applied the Cliff $\delta$ [4], to measure the independent effect and interpreted the effect size as follows [34]: $|\delta|<0.147$ as negligible, $0.147\leq|\delta|<0.33$ as small, $0.3\leq|\delta|<0.474$ as medium, and $0.474\leq|\delta|$ as large.
@@
We categorized the patterns that emerged from the RDD analysis based on two key characteristics: (1) trend and (2) slope. The trend characteristic captures whether the code churn exhibits an upward or downward trend when comparing pre- and post-GenAI adoption periods. The slope characteristic captures whether and how the slope of the trend line changes before and after GenAI adoption. The Ordinary Least Squares (OLS) model used as part of RDD requires a minimum of five weeks of data to estimate the four parameters intercept, time trend, treatment effect, and interaction, while maintaining positive degrees of freedom [17] to ensure that there are enough data points to estimate the model parameters without overfitting. After applying a threshold of at least one commit per week over five weeks, we had to exclude two repositories without sufficient data in the pre-GenAI adoption period.
For the 149 repositories included, we identified four patterns:
@@
These patterns provide a useful framework for analyzing how code churn metrics change after GenAI adoption in different project contexts. To assess the robustness and causal interpretability of the RDD-based patterns, we conducted robustness checks. First, in addition to each repository’s project-specific adoption point (i.e., the first commit explicitly reporting GenAI assistance), we estimated parallel RDD models using two global cutoff dates corresponding to major GenAI releases: GitHub Copilot (June 2021) and ChatGPT (November 2022). Second, we assessed robustness to bandwidth selection by estimating the RDD under multiple symmetric temporal windows around the cutoff (±90, ±180, and ±360 days), while keeping all other model specifications constant. Third, we conducted placebo tests by shifting each repository’s cutoff date forward and backward in time and re-estimating the same models. Across all robustness checks, the results remained stable, indicating that our main RDD findings do not depend on arbitrary cutoff choices or temporal window specifications.
Finally, to examine heterogeneity across different forms of GenAI usage, we further disaggregated the RDD analysis by the GenAI-assisted tasks identified in RQ1 (e.g., generation, optimization, and maintenance). We additionally performed a manual verification of every subcode assigned to each task in Table II, including those subcodes likely to influence code churn. The full set of GenAI generation task categories includes code, test data, comment, test file, regex, test method, Zod schema, test class, and configuration. GenAI optimization task categories encompass all subcodes of code refactoring and code improvement. GenAI maintenance task categories include comment revision, color suggestions, dependency upgrades, and version updates.
Table IX: Significant RDD churn discontinuities by task category (Generation/Optimization/Maintenance).
Churn Type Task #Significant Sum sig. Not sig.
File-based Generation 10 $|$ 6 16 68
Optimization 1 $|$ 1 2 12
Maintenance 0 $|$ 1 1 2
Line-based Generation 9 $|$ 6 15 69
Optimization 2 $|$ 2 4 10
Maintenance 0 $|$ 2 2 1
@@ 5.2 Results
Table VII illustrates the variations in code churn of the studied repositories. Of the 151 repositories with self-admitted GenAI usage, 119 had a significant difference in file-based churn, and 120 had a significant difference in line-based churn ($p<0.05$). Eleven repositories had an increasing file-based churn with a medium-to-large effect size, and 15 had an increasing line-based churn with a medium-to-large effect size. A decreasing churn was more common: 49 repositories had a decreasing file-based churn with a medium-to-large effect size, and 49 repositories had a decreasing line-based churn with a medium-to-large effect size.
@@
These results are contrary to our expectations because the GitClear report was very bold in claiming that code churn increased for the projects they studied, suggesting a downward pressure on code quality [10]. While we observed that some repositories have an increasing trend in code churn, both the overall trend and the trend and many individual repositories point to a decreasing code churn over time. Therefore, with our data and methodology, we cannot confirm this claim.
Table VIII summarizes the results of our RDD analysis. We observed that only 26 (file-based) respectively 30 (line-based) repositories showed significant code churn trends ($p<0.05$). For file-based churn, an overall upward trend with a negative slope after the cutoff date was most common (12 repositories). For line-based churn, an overall downward trend with a negative slope was most common (12 repositories). However, there were almost as many repositories (10) with an overall downward trend, but a positive slope after the cutoff date. Figure 2 presents examples of all four patterns that we observed, and the complete RDD results are available as part of our replication package. In summary, we found 15 repositories with a significant upward trend in file-based churn and 15 with a significant upward trend in line-based churn. However, most of them had a negative slope. In addition, 11 projects had a significant downward trend in file-based churn and 15 had a significant downward trend in line-based churn. Overall, we cannot conclude that there is a general trend toward increasing code churn.
To deepen our understanding of which GenAI-assisted tasks (RQ1) are most strongly associated with churn changes, we have further disaggregated the RDD results by task (see Table IX). Across all models, significant discontinuities are concentrated in the generation category, including the generation of code, test data, test methods, regular expressions, and comments. In contrast, optimization and maintenance tasks do not show consistent patterns. In summary, this task-level analysis reveals that GenAI-assisted generation tasks appear more rework-prone than other forms of assistance.
Summary RQ3: Our results revealed that for most of the repositories analyzed, there was no significant change in code churn after GenAI adoption. We did find 15 repositories with an overall upward trend in line-based code churn after the first GenAI mention. However, for most of them, the slope was negative. In addition, we also found 15 repositories with a downward trend. Among the GenAI tasks identified in RQ1, generation tasks show a stronger impact on code churn than other tasks. These results indicate that more research is required to understand why certain projects or GenAI tasks are affected and others are not, and how higher (or lower) code churn relates to the long-term maintainability of software projects.
@@ 6 Discussion
In this section, we discuss and contextualize the results of our three research questions and summarize the implications for software developers and researchers.
@@ 6.1 RQ1: Reasons for Mentioning GenAI Tools
Our taxonomy provides a multi-dimensional characterization (task, content type, purpose), where prior studies only categorize tasks assigned to GenAI tools. Our analysis revealed that developers primarily use GenAI tools for code generation, natural language translation, and code refactoring. Tufano et al. [43] explored mentions of ChatGPT in commits, PRs, and issues. They identified that the three most common task categories were feature implementation and enhancement, software quality, and documentation. In our study, we present a more fine-grained and comprehensive categorization of tasks automated by both ChatGPT and GitHub Copilot. For studies targeting GitHub repositories, it is crucial to consider GitHub Copilot as well, because (1) unlike ChatGPT, it is a tool tailored to software development, and (2) it is more deeply embedded in developers’ workflows (their local editors, but also into the GitHub platform as a whole). Moreover, we complement the task categories by specifically discussing content and usage purposes. In addition, we identified patterns of human intervention. Hou et al. [16] reviewed literature on LLMs for software engineering. They found that software engineering research has a strong focus on code generation and program repair. We complement this observation with a detailed taxonomy of how open-source developers use LLM-based tools in their projects. In addition to code generation, we found that internationalization and natural language translation are common use cases for LLMs in open-source software projects (49 instances, see Table II). This finding is aligned with Tufano et al. [43], who identified 12 instances of ChatGPT usage for internationalization. Such use cases highlight the need to support not only programming tasks but also broader software development activities. We further found instances of projects regulating the usage of GenAI tools, which we analyzed in more detail as part of RQ2. While our study partially confirms previous studies on software development tasks being automated using GenAI tools, we contribute three novel perspectives: (1) some developers deeply care about acknowledging GenAI usage in open-source software, (2) open-source maintainers try to actively guide and regulate GenAI usage, and (3) issues with generated code can trigger human interventions in open-source software projects.
@@
Based on our findings, we recommend researchers to investigate developers’ rationale behind self-admitted (vs. hidden) GenAI usage. Our notion of self-admitted GenAI usage, inspired by self-admitted technical debt [30], can be a valuable lens for studying GenAI usage in practice. Of course, only a fraction of the generated software artifacts contain GenAI mentions, and the artifacts that are documented might not be representative of the overall GenAI usage. Better understanding when and why developers decide to self-admit GenAI usage is one potential direction for future work. Moreover, our annotated dataset of self-admitted GenAI usage can serve as a starting point to build a tool to automatically identify true positive GenAI mentions according to the definition presented in Section 2.2. An improved and scaled detection of self-admitted GenAI usage would allow researchers to build larger datasets that could then enable more comprehensive studies on code quality and maintainability of generated code.
Software developers can browse our taxonomy of tasks, content types, and purposes to identify potential applications of GenAI tools in their projects. One central aspect is whether to establish guidelines clarifying in which cases project maintainers require contributors to disclose and acknowledge GenAI usage (see also Section 6.2).
@@
We found that some acknowledgments were combined with warnings about the potential negative implications of GenAI tool usage. Sometimes, GenAI tools were also blamed for issues. However, acknowledgment can also serve a positive purpose, e.g., documenting prompts. The question arises not only when to acknowledge GenAI usage, but also which context to document beyond the tool name (which we focused on). In which cases does it make sense to document complete prompts and where and how should one document the generation context? Such questions can be addressed both from a scientific and from a practical perspective. A more standardized approach for documenting GenAI contributions is required, since most self-admitted GenAI usages did not document the generation context beyond brief summaries. As discussed above, we found cases documenting prompts behind GenAI usage. GenAI tool builders (e.g., of IDE plugins) could offer an option to record these contexts, enabling transparent acknowledgment of GenAI usage and prompt reuse.
@@ 6.2 RQ2: Existing Guidelines for GenAI Usage
Motivated by the purpose categories Documentation and Acknowledgement and Guidance and Best Practices that we identified while answering RQ1 (see Table V), we further explored the policies and usage guidelines for GenAI tools that we found (see Table VI). Their content ranged from encouraging developers to use GenAI tools to prohibiting their usage entirely.
@@
Our developer survey confirmed the broad spectrum of positions that cover ethical, legal, and practical considerations. Mentioned aspects include the unclear copyright situation of the training data, the unclear implications for generated content, data privacy risks when sharing inputs with GenAI systems, and concerns regarding code quality and maintainability. Moreover, a majority of our survey participants agreed that the regulation of GenAI usage is necessary in open-source projects. In relation to that, participants argued for transparent disclosure of GenAI usage and also for documenting the generation context. It is unclear how much transparency is required and what purposes it can serve: Is a binary flag sufficient? Or is it better to document the percentage of generated content, as suggested by a participant? Or the whole prompt? Do only manually written prompts need to be disclosed, or also system prompts? This aspect is aligned with the questions raised in the discussion for RQ1 about prompt context.
@@
Our results suggest that software developers, especially those maintaining open-source software projects, should articulate a clear position regarding GenAI usage in their projects. The spectrum of possible positions ranges from a general recommendation to use GenAI tools, over recommendations for specific tools and use cases, to more restrictive policies requiring an extensive peer review of generated content, or policies prohibiting GenAI usage completely. Open-source projects should clearly communicate expectations regarding GenAI usage to their contributors. For downstream consumers of open-source dependencies, explicit GenAI policies serve as a signal of due diligence that may influence their dependency selection.
@@
Our analysis of policies, guidelines, and developers’ positions regarding GenAI regulation provides a solid foundation for researchers to design and conduct further studies on how software projects regulate GenAI usage and how such regulations impact development activity. An idea worth exploring is whether existing GenAI tools could be augmented to capture provenance information during generation that could be automatically added to source code comments, comment messages, or other artifacts such as Software Bills of Materials (SBOMs) [33] or Software Bill of Materials for AI (SBOM for AI) [51]. There are already open-source projects that extensively document prompts in commit messages.111github.com/cloudflare/workers-oauth-provider/commit/adcb… This provenance information is essential to study the long-term impact of code generation on maintainability, but also to facilitate software supply chain transparency and vulnerability management. Researchers can contribute to the development of standardized metadata formats to capture provenance and traceability information of code and other software artifacts.
@@ 6.3 RQ3: Impact of GenAI Usage on Code Churn
Our results for RQ3 challenge popular narratives about the impact of GenAI on software development. Contrary to claims in the GitClear report, which was extensively discussed in the software development community [15, 32], we did not find an increasing code churn after GenAI adoption. The overall trend we observed pointed in the opposite direction, i.e., we noticed a decreasing average code churn. This is in line with a study by Grewal et al. [13] which examined how ChatGPT-generated code is integrated into GitHub projects. They found that approximately 54% of the generated code lines were integrated and only 2.5% of them were later modified. However, our RDD analysis revealed that three repositories (3/26, 11.5%) had a significant upward file-based code churn trend with a positive slope ($p<0.05$). This indicates that a subset of repositories exhibited a progressively faster increase in code churn after GenAI adoption. To contextualize this finding, we note that a larger proportion of repositories (12/26, 46.2%) showed an upward trend with a negative slope, i.e., churn increased, but the increase decelerated over time. In addition, 26.9% (7/26) showed downward trends. This heterogeneity suggests that GenAI’s impact on code churn is context-dependent rather than uniformly negative, as suggested by the GitClear report.
Building on our RQ3 results, we recommend researchers to explore the factors that contribute to increased code churn. The patterns we identified using our RDD analysis are a valuable lens for clustering projects. This clustering can inform a detailed qualitative study of projects that exhibit similar patterns. The difference between our results and the GitClear report can be partially attributed to the methodological differences between the studies. While GitClear used a global cutoff date, we used the first GenAI mention in a repository as a proxy for GenAI adoption, thus following a more fine-grained approach. Moreover, we introduced file-level and line-level code churn and calculated churn at the project-level and globally. Our definitions and the code in our replication package enable other researchers to consider code churn in their own studies.
For software developers, we further suggest to monitor the impact of GenAI on the software projects they maintain or contribute to. Our results suggest that the impact of GenAI adoption on the development activity in software projects might not be as clear as suggested by the GitClear report. Considering that we did notice an increasing code churn in several projects, it is nevertheless important for project maintainers to monitor the development activity and the quality of contributions. Going forward, we might extend our implementation to calculate code churn into a tool that project maintainers can easily integrate into their repositories.